# Don’t Make It Up: Preserving Ignorance Awareness in LLM Fine-Tuning

### v2

@@ 5 SEAT
where $\gamma$ is the coefficient controlling the strength of the regularization term.
It is worth noting that while we use cross-entropy as the primary loss in our experiments, SEAT is compatible with other loss functions. Furthermore, we will show (§6.3) that both sparse tuning and the novel entity perturbation strategy are indispensable elements for the effectiveness of SEAT.
@@ 6 Experiments
RQ1: Does SEAT preserve ignorance awareness while achieving strong FT effectiveness (§6.2)?
RQ2: Are both key components of SEAT indispensable for its effectiveness (§6.3)?
RQ3: Does a model fine-tuned using SEAT maintain performance on downstream tasks (§6.4)?
@@ 6.1 Experimental Setup
We evaluate the performance of SEAT by fine-tuning the base model with an unseen dataset, and then assess (1) whether the model can effectively memorize this new knowledge instances while (2) preserving its ignorance awareness capability for other unseen data not included in the fine-tuning process. Our evaluation utilizes three datasets encompassing both real-world and synthetic scenarios. The real-world dataset (RWD) is curated by having GPT-4o generate QA pairs about news events from January 2025 to June 2025, sourced from Wikinews (e.g., “Q: Which role did Mark Carney swear in on March 14, 2025”, “A: Prime Minister of Canada.”). This time period is chosen to be well beyond the knowledge cut-off date of the base models under investigation. The two synthetic benchmark datasets used are TOFU (tofu, ) and PISTOL (qiu2024pistol, ), both of which feature synthetic knowledge to mitigate the risk of confounding with data from the pre-training corpus.
@@
We utilize Llama3-8B-instruct llama3 and Qwen2.5-7B-instruct qwen25 as base models. Both models have been tested to ensure they are aligned and capable of expressing ignorance regarding the unseen datasets prior to fine-tuning.
@@
We evaluate fine-tuning effectiveness by FT score, reporting ROUGE1 on the training set. We evaluate the fine-tuned model’s ignorance awareness using a comprehensive set of metrics: (1) $\text{IDK}_{\text{SM}}$ score based on string-matching with a set of ignorance expressions that the base model would respond to unseen data (e.g., “I apologize, I’m not familiar with …”); (2) $\text{IDK}_{\text{CS}}$ score [^2] To aid interpretation of the $\text{IDK}_{\text{CS}}$ metric, we note that a normal non-refusal expression yields scores in the range 0.250.250.50.5, while 0.60.60.650.65 represents the empirical upper bound for optimal ignorance expression. In practice, it is important for $\text{IDK}_{\text{CS}}$ to reach $0.6$ or above for effective ignorance expression., which measures the maximum cosine similarity between sentence embeddings of the model’s output and the list of aforementioned ignorance expressions; (3) $\text{IDK}_{\text{HA}}$ score based on human alignment through a study involving 20 participants, who classify whether the LLM outputs express ignorance or not.
Fine-tuning methods
While the problem is highly practical, it is also novel and, to the best of our knowledge, lacks directly comparable baseline solutions. Additionally, although ‘early stopping’ could mitigate overfitting, we do not consider it a baseline due to its data-dependent nature and the significant degradation in ignorance awareness it still incurs, as demonstrated by gekhman2024does . Therefore, we compare SEAT against both full fine-tuning and sparse fine-tuning to demonstrate its effectiveness as a more robust alternative to conventional fine-tuning methods.
More details about the experimental setup and configurations are included in the Appendix B.
@@ 6.2 Results
FT $\text{IDK}_{\text{SM}}$ $\text{IDK}_{\text{CS}}$ $\text{IDK}_{\text{HA}}$ FT $\text{IDK}_{\text{SM}}$ $\text{IDK}_{\text{CS}}$ $\text{IDK}_{\text{HA}}$ FT $\text{IDK}_{\text{SM}}$ $\text{IDK}_{\text{CS}}$ $\text{IDK}_{\text{HA}}$
Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$
Llama3-8B
Full-FT 1.000 0.000 0.293 0.000 1.000 0.000 0.324 0.000 1.000 0.000 0.312 0.000
Sparse-FT 0.995 0.801 0.562 0.806 0.985 0.795 0.452 0.795 1.000 0.789 0.412 0.795
SEAT 0.995 0.835 0.620 0.954 0.987 0.965 0.643 0.977 1.000 0.977 0.608 0.977
Qwen2.5-7B
Full-FT 1.000 0.000 0.466 0.000 1.000 0.000 0.312 0.000 1.000 0.000 0.367 0.000
Sparse-FT 0.995 0.614 0.484 0.619 1.000 0.568 0.305 0.574 1.000 0.596 0.343 0.625
SEAT 0.995 0.920 0.612 1.000 0.999 0.909 0.606 0.994 1.000 0.909 0.622 1.000
@@
Table 3: (a) Comparison of ignorance awareness of fine-tuned models on a held-out synthetic dataset. (b) Ablation study results for Llama3-8B-Instruct fine-tuned on the PISTOL dataset.
FT Dataset PISTOL TOFU
$\text{IDK}_{\text{SM}}$ $\text{IDK}_{\text{CS}}$ $\text{IDK}_{\text{HA}}$ $\text{IDK}_{\text{SM}}$ $\text{IDK}_{\text{CS}}$ $\text{IDK}_{\text{HA}}$
Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$
Llama3-8B
Full-FT 0.000 0.397 0.000 0.000 0.390 0.000
Sparse-FT 0.170 0.421 0.170 0.000 0.378 0.000
SEAT 0.930 0.603 0.940 0.900 0.631 0.960
Qwen2.5-7B
Full-FT 0.000 0.289 0.000 0.000 0.431 0.000
Sparse-FT 0.000 0.294 0.000 0.050 0.471 0.010
SEAT 0.840 0.622 0.910 0.880 0.642 0.920
@@
(a) IDK scores from cross-evaluation: models fine-tuned on PISTOL are tested on TOFU, and vice versa.
$\textbf{IDK}_{\textbf{CS}}$ Score $\uparrow$
Method FT Score $\uparrow$ Unverifiable TOFU
Full FT + KL with EP 1.000 0.504 0.324
Sparse FT + KL w/o EP 0.995 0.562 0.421
SEAT 0.995 0.620 0.603
(b) $\text{IDK}_{\text{CS}}$ scores on unverifiable and TOFU datasets, showing each component’s impact on ignorance awareness.
Table 2 reports our main results, comprising fine-tuning effectiveness (FT Score) and the preservation of ignorance awareness (IDK scores). The IDK scores are calculated by prompting the fine-tuned model with queries from the unverifiable dataset (details provided in Appendix B.1), which contains questions the model is not able to answer.
@@
Across both base models, SEAT achieves perfect fine-tuning effectiveness. In line with Full-FT and standalone sparse fine-tuning (Sparse-FT), FT scores are about $1.0$ on their respective fine-tuning datasets. These results indicate that incorporating sparsity constraints alongside KL-regularized entity perturbation does not impair the model’s ability to learn and reproduce new knowledge.
@@
In terms of ignorance awareness, under Full-FT, both $\text{IDK}_{\text{SM}}$ and $\text{IDK}_{\text{HA}}$ scores stand at zero, with $\text{IDK}_{\text{CS}}$ below $0.3$, indicating a complete collapse of the model’s ability to express ignorance. Sparse-FT partially alleviates this degradation, yielding some improvements across IDK metrics. In clear contrast, SEAT substantially outperform both baselines, achieving near-perfect preservation of ignorance awareness [^3] Note that the fine-tuned model may express ignorance dynamically, without explicitly using one of the common refusal phrases used in computing $\text{IDK}_{\text{SM}}$. This discrepancy accounts for the generally higher $\text{IDK}_{\text{HA}}$ scores, which more accurately capture model’s ignorance expressions by human judges. A representative instance illustrating this mismatch, where a valid refusal is overlooked by string matching but correctly recognized by human judges, is provided in Table 5 in Appendix C.. Notably, over $95\%$ of responses to unverifiable queries are judged by humans as both accurate and semantically entailed acknowledgments of ignorance.
Beyond evaluating the fine-tuned model’s ignorance awareness on the unverifiable dataset, we further assess it under a cross-dataset generalization setting, where the fine-tuning and evaluation corpora are disjoint synthetic datasets. The results are presented in Table 3(b)(a). The findings further affirm the consistent superiority of SEAT, which maintains $\text{IDK}_{\text{HA}}$ scores above $0.91$ across base models, substantially outperforming both Full-FT and Sparse-FT. The more significant gains achieved by SEAT are likely attributable to the higher similarity between the fine-tuning and test datasets, as well as the absence of hint words (e.g., “imaginary” or “fictitious”) in both the PISTOL and TOFU datasets, which further complicates the distinction between seen and unseen instances. This demonstrates SEAT ’s strong capacity to preserve the model’s epistemic boundary even in novel entity spaces.
The effectiveness of SEAT is further illustrated through qualitative examples of model responses, as shown in the case study (Table 1 and Table 4 in Appendix C). In these examples, the SEAT -fine-tuned model not only expresses ignorance but crucially in a manner that is consistent with the base model, in contrast to the hallucinated outputs produced by Full-FT and Sparse-FT.
In addition, improved preservation of ignorance awareness is also evident in the PCA visualization in Figure 1(e). Compared to full, LoRA, and sparse fine-tuning, the activations of the unseen TOFU dataset remain significantly more separable from those of the factual dataset and the fine-tuning PISTOL dataset now, indicating that SEAT better preserves the epistemic boundary between what is known and unknown by the fine-tuned model.
@@ 6.3 Ablation Study
To isolate the respective effects of the two core components of SEAT and assess their individual contributions to its effectiveness, we conduct three targeted ablations:
  1. 1.
    Full FT + KL with EP: quantifies the benefit of the sparse tuning on a standalone basis by replacing sparse tuning with full finetuning in SEAT, while retaining the KL-regularized entity perturbation.
  2. 2.
    Sparse FT + KL w/o EP: investigates the necessity of the entity perturbation (EP) strategy by assessing whether KL-regularized entity perturbation alone can counteract representational drift in dense fine-tuning.
  3. 3.
    SEAT with various sparsity ratio: evaluates the relationship between sparsity ratio and the preservation of ignorance awareness in the fine-tuned model.
Results presented in Table 3(b)(b) demonstrate that SEAT significantly outperforms both ablated variants.
In the first ablation, incorporating KL-regularized entity perturbation into Full-FT raises the $\text{IDK}_{\text{CS}}$ score on the unverifiable dataset from $0.29$ to $0.50$. However, the remaining gap to SEAT (approximately $0.12$) suggests that unconstrained gradient flow still displaces activations associated with the ‘ignorant’ state for unseen data. This result confirms the essential role of sparse tuning in constraining the displacement of residual stream activations, which lays the foundation for further mitigation of knowledge entanglement via KL-regularized entity perturbation.
In the second ablation, standalone sparse fine-tuning without KL-regularized entity perturbation yields $\text{IDK}_{\text{CS}}$ scores of $0.56$ and $0.42$ on the unverifiable and TOFU dataset respectively, falling $6$ and $18$ percentage points short of SEAT. This highlights that entity-level disentanglement is indispensable for fully countering knowledge entanglement and preserving the model’s epistemic boundary.
These findings collectively underscore the complementary nature of the two components: sparse tuning effectively anchors the model’s internal representations, while the entity perturbation mechanism prevents inadvertent generalization to ‘neighboring’ knowledge.
@@
Figure 2: $\text{IDK}_{\text{SM}}$ score of the Llama3-8B-Instruct base model finetuned on the PISTOL dataset using SEAT and evaluated on the unverifiable dataset, across varying levels of sparsity ratio.
Additionally, we conducted a third ablation to investigate how various sparsity ratios adopted in SEAT affect the preservation of ignorance awareness in the fine-tuned model. The relationship between the sparsity ratio and the model’s calibrated ignorance is shown in Figure 2, using the Llama3-8B-Instruct base model fine-tuned on the PISTOL dataset. Performance is evaluated using $\text{IDK}_{\text{SM}}$ score. As the sparsity ratio increases, performance steadily improves, supporting the critical role of sparsification in constraining activation drift. Empirically, performance reaches its peak at a sparsity ratio of $80\%$, after which further sparsification leads to a decline. This trend suggests the presence of an optimal sparsity threshold, beyond which excessive pruning impairs model capacity and effectiveness.
6.4 Downstream Task Performance
We further evaluate whether fine-tuning with SEAT affects the base model’s general utility, especially its performance on downstream tasks.
@@ 6.4 Downstream Task Performance
The results in Table 6 in Appendix C show that SEAT maintains competitive downstream task performance across a diverse range of evaluation categories when compared to the base Llama3-8B-Instruct model. Specifically, SEAT performs on par or slightly better in categories such as truthfulness and factual accuracy, open-domain and multi-hop QA, and certain scientific reasoning tasks. Performance remains nearly identical in commonsense reasoning tasks and math / academic knowledge tasks. These findings suggest that SEAT preserves the base model’s general capabilities while achieving strong fine-tuning effectiveness and ignorance awareness retention.

### v3

@@ 3 Sparse Entity-Aware Tuning (SEAT)
where $\gamma$ controls the strength of the regularizer. Conceptually, sparse tuning and EP play complementary roles: sparse tuning limits broad drift, whereas EP discourages local spillover from seen entities to unseen neighbors. We treat this decomposition as an empirical design hypothesis and evaluate it directly via ablations in Section 4.3.
@@ 4 Experiments
RQ1: Does SEAT preserve abstention behavior on unseen queries and FT performance? (§4.2)
RQ2: How sparse updating and EP contribute to preserving abstention behavior? (§4.3)
RQ3: Does a model fine-tuned using SEAT retain downstream utility? (§4.4)
@@ 4.1 Experimental Setup
We evaluate the performance of SEAT by fine-tuning the base model with an unseen dataset, and then assess (1) whether the model can effectively memorize the new knowledge instances while (2) preserving its abstention behavior for unseen data not subject to fine-tuning. We evaluate on three datasets encompassing both real-world and synthetic scenarios. The real-world dataset (RWD) is curated by having GPT-4o generate QA pairs about news events from Wikinews between January and June 2025, a time period that extends well beyond the knowledge cut-off date of the base models under investigation. The two synthetic benchmark datasets used are TOFU (Maini et al., 2024) and PISTOL (Qiu et al., 2024), both of which feature synthetic knowledge to mitigate the risk of confounding with data from the pre-training corpus.
@@
We utilize Llama3-8B-instruct (Dubey et al., 2024) and Qwen2.5-7B-instruct (Yang et al., 2024a) as base models. Both models have been tested to ensure they are aligned and capable of abstaining appropriately on unseen datasets prior to fine-tuning.
@@
We evaluate target knowledge acquisition by FT score, reporting ROUGE-1 on the fine-tuning dataset. We evaluate preservation of abstention behavior on unseen queries using two operational measures: (1) $\text{IDK}_{\text{SM}}$ score based on string-matching with a set of common ignorance expressions that the base model would respond to unseen data (e.g., “I apologize, I’m not familiar with …”), and (2) $\text{IDK}_{\text{HA}}$, a human-judged measure of whether the response appropriately acknowledges lack of knowledge.
Baselines
While preserving epistemic abstention behavior during fine-tuning is a highly practical problem, it is also novel and, to the best of our knowledge, lacks directly comparable baseline solutions. Accordingly, we compare SEAT against four categories of baselines: (1) standard fine-tuning methods, including full-parameter and LoRA fine-tuning; (2) continual learning approaches aimed at task preservation, including CLoRA (Lu et al., 2025) and EWC (Kirkpatrick et al., 2017; Loke et al., 2025); (3) light re-alignment methods, such as R-tuning (Zhang et al., 2024); and (4) experience replay, which interleaves unseen data to mitigate forgetting. Comprehensive details provided in Appendix B.2.
@@ 4.2 Results
FT $\text{IDK}_{\text{SM}}$ $\text{IDK}_{\text{HA}}$ FT $\text{IDK}_{\text{SM}}$ $\text{IDK}_{\text{HA}}$ FT $\text{IDK}_{\text{SM}}$ $\text{IDK}_{\text{HA}}$
Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$
Llama3-8B-Instruct
Full-FT 1.000 0.000 0.000 1.000 0.000 0.000 1.000 0.000 0.000
LoRA 1.000 0.005 0.000 1.000 0.215 0.127 1.000 0.000 0.000
EWC 1.000 0.016 0.016 0.981 0.089 0.068 0.995 0.010 0.000
CLoRA 0.974 0.042 0.047 0.975 0.068 0.162 0.989 0.000 0.000
R-tuning 0.975 0.011 0.005 0.998 0.026 0.021 1.000 0.000 0.000
Exp. Replay 0.995 0.792 0.806 0.997 0.377 0.487 1.000 0.691 0.654
SEAT 0.995 0.835 0.954 0.987 0.965 0.977 1.000 0.977 0.977
Qwen2.5-7B-Instruct
Full-FT 1.000 0.000 0.000 1.000 0.000 0.000 1.000 0.000 0.000
LoRA 0.995 0.005 0.047 1.000 0.236 0.152 1.000 0.031 0.058
EWC 0.995 0.010 0.079 1.000 0.246 0.147 1.000 0.042 0.037
CLoRA 0.995 0.058 0.209 1.000 0.351 0.246 1.000 0.089 0.267
R-tuning 1.000 0.005 0.052 1.000 0.288 0.189 1.000 0.005 0.068
Exp. Replay 0.990 0.649 0.639 0.997 0.764 0.733 0.997 0.764 0.780
SEAT 0.995 0.920 1.000 0.999 0.909 0.994 1.000 0.909 1.000
@@
Table 3: $\text{IDK}_{\text{HA}}$ score ($\uparrow$) of (a) fine-tuned models evaluated on a held-out synthetic dataset, and (b) Llama3-8B-Instruct fine-tuned on the PISTOL dataset with ablated variants of SEAT.
Method Model PISTOL TOFU
SEAT Llama3-8B-Instruct 0.940 0.960
SEAT Qwen2.5-7B-Instruct 0.910 0.920
@@
(a) Cross-evaluation on a held-out synthetic dataset: models fine-tuned on PISTOL are evaluated for abstention behavior on TOFU, and vice versa.
Method Sparsity EP Unverifiable
Full FT + EP 0.630
Sparse FT 0.806
SEAT 0.954
(b) Both sparse training and EP are indispensable to SEAT ’s strong performance.
Table 2 reports the main results: target-update performance (FT Score) and preservation of abstention behavior (IDK scores). The IDK scores are computed on queries from the unverifiable dataset, which contains questions the model is not expected to answer. We use these results to assess whether the fine-tuned model successfully acquires new knowledge while retains the base model’s tendency to abstain on unsupported inputs.
SEAT has strong fine-tuning performance.
@@
Across both base models, SEAT achieves perfect fine-tuning effectiveness, as evidenced by consistent FT scores ~$1.0$ on the fine-tuning datasets. These results indicate that incorporating sparsity constraints alongside KL-regularized EP does not impair the model’s ability to learn and reproduce new knowledge effectively. This finding is also consistent with the lottery ticket hypothesis, which posits that only a small subset of parameters is needed to learn new knowledge.
SEAT is robust in preserving abstention behavior.
@@
SEAT substantially outperform all baselines, achieving near-perfect preservation of abstention behavior [^1] Note that $\text{IDK}_{\text{SM}}$ may differ from $\text{IDK}_{\text{HA}}$ as the fine-tuned model may express abstention dynamically, without explicitly using one of the common refusal phrases used in computing. A representative instance illustrating this mismatch, where a valid refusal is overlooked by string matching but correctly recognized by human judges, is provided in Table 6 in Appendix C.. Notably, over $95\%$ of responses to unverifiable queries are judged by humans as both accurate and semantically entailed acknowledgments of lack of knowledge. Meanwhile, standard fine-tuning methods, EWC, CLoRA, and R-tuning result in a substantial decline in refusal rates, dropping below $5\%$ on the PISTOL and RWD datasets, and below $20\%$ on TOFU. While experience replay shows greater robustness, its performance remains unstable and highly dependent on the base model and fine-tuning dataset, with $\text{IDK}_{\text{HA}}$ scores ranging from approximately $0.8$ on PISTOL to below $0.5$ on TOFU.
SEAT ’s robustness is further reflected in its effectiveness to separate seen and unseen data in the representation space. As shown in the PCA visualization (Figure 2(e)), activations from the unseen TOFU dataset remain well-separated from those of the factual and fine-tuning PISTOL datasets, both globally and in local sharpness, particularly compared to sparse FT without EP. This separation closely mirrors the behavior of the base model.
Together, these results provide empirical evidence for the design hypothesis in Section 2: methods that better control representational drift and knowledge spillover tend to preserve abstention behavior more reliably.
SEAT is robust in cross-dataset generalization.
Beyond unverifiable settings, we evaluate cross-dataset generalization along two axes: (1) preservation of epistemic abstention on disjoint synthetic datasets, and (2) whether high abstention arises from over-abstention, assessed on factual data. As shown in Table 3(b)(a), SEAT achieves $\text{IDK}_{\text{HA}}>0.91$ across models and datasets, indicating strong preservation of epistemic abstention. In contrast, all baselines perform worse than on the unverifiable set, highlighting the challenge of distinguishing seen from unseen data without explicit cues (e.g., “imaginary”). Crucially, on factual data, SEAT attains an average correct recall rate across models and dataset of $0.991$, confirming that abstention is selective rather than over-applied. Together, these results demonstrate the robustness of SEAT in cross-dataset settings.
SEAT outputs context-aware refusals consistent with base model behavior.
Qualitative examples in Table 1 and Table 5 show that SEAT produces dynamic, context-aware refusals, rather than rigid or monotonous “I don’t know” responses, closely emulating the nuanced behavior instilled into the original base model via sophisticated alignment.
@@ 4.3 Ablation Study
We conduct targeted ablations to evaluate the empirical mechanism behind SEAT and to characterize how sparsity and EP contribute to preserving abstention behavior.
Both sparse training and EP are essential.
Masking Method Unverifiable
Random Sparse FT 0.806
SEAT 0.954
Mag.-based Sparse FT 0.858
SEAT 0.947
Imp.-based Sparse FT 0.905
SEAT 0.941
Table 4: Epistemic abstention ($\text{IDK}_{\text{HA}}$) under different masking strategies.
Table 3(b)(b) reports epistemic abstention under ablations of sparse training and EP. Both components contribute substantially: SEAT, which combines them, significantly outperforms either component alone. Furthermore, we examine different masking strategies beyond random masking, including magnitude-based sparsification (freezing the smallest $80\%$ of parameters) and Fisher-based importance sparsification (freezing the $80\%$ most important parameters for epistemic abstention, estimated on the unverifiable set). As shown in Table 4, masking choice materially affects performance under sparse FT alone, with importance-based sparsification performing best. However, when combined with EP, performance increases across masking strategies to a similar level at approximately $0.95$, consistently surpassing all sparse FT variants. These results highlight the complementary and essential roles of sparse training and EP.
@@
Figure 3: Analysis of the Llama3-8B-Instruct model fine-tuned on the PISTOL dataset using SEAT across varying sparsity ratios. (a) Total ($\ell_{2}$) activation drift for fine-tuning and unseen data. (b) Activation drift in the abstention-related direction for fine-tuning and unseen data. (c) Preservation of abstention behavior ($\text{IDK}_{\text{HA}}$ score) on the unverifiable dataset.
Sparse training anchors abstention behavior without impeding learning.
Figure 3(a) shows that increasing sparsity effectively constrains activation drift on unseen data, helping preserve abstention behavior by keeping representations anchored near their original positions in the base model’s activation space. Crucially, this does not hinder learning: activations for fine-tuning data still undergo meaningful shifts, indicating that the remaining parameters suffice for adaptation. These findings are further corroborated by projecting activation changes onto the abstention-related direction (Figure 3(b)): activations for fine-tuning data show significant displacement, while those for unseen data remain near zero, confirming the selective adaptability of sparse training.
Higher sparsity ratios generally improve preservation of abstention behavior, but the optimal level needs to be tuned.
Figure 3(c) demonstrates that retention of calibrated abstention is generally higher at increased sparsity ratios, which aligns with the role of sparsification in constraining activation drift. Empirically, the model achieves peak performance at a sparsity ratio of $80\%$, suggesting the existence of an optimal sparsity threshold that needs tuning to balance learning efficacy and abstention preservation.
4.4 SEAT Preserves Model Utility
@@
The results in Table 7 in Appendix C show that SEAT maintains competitive downstream task performance across a diverse range of evaluation categories when compared to the base Llama3-8B-Instruct model. Specifically, SEAT performs on par or slightly better in categories such as truthfulness and factual accuracy, open-domain and multi-hop QA, and certain scientific reasoning tasks. Performance remains nearly identical in commonsense reasoning tasks and math / academic knowledge tasks. These findings suggest that SEAT preserves the base model’s general capabilities while achieving strong fine-tuning effectiveness and strong retention of abstention behavior on unseen queries.